OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#7427 — Lille Switchs Update
Scheduled Maintenance Report for Web Cloud
Completed
We will update switches managing Lille's xDSL infrastructure.

Update(s):

Date: 2012-10-09 19:22:16 UTC
An inconsistency of the configuration was at the origin of the problem. The mac addresses will not synchonize anymore between the Lille routers. The problem is currently fixed.

Date: 2012-10-09 10:47:37 UTC
Some clients lose internet connection.
It is random and does not affect all customers.

We seek the origin of the problem.

.

Date: 2012-10-09 06:13:00 UTC
Both switchs were updated.

5 minutes cut-off is expected.

Date: 2012-10-09 06:11:55 UTC
Done.

Moving to the second.

Date: 2012-10-09 06:11:36 UTC
We will start with the first switch.
Posted Oct 09, 2012 - 06:11 UTC